# Bayesian Optimization with Preference Exploration using a Monotonic Neural Network Ensemble

### v2

@@ 6 Conclusion
This paper introduces BOPE-MoNNE (Bayesian Optimization with Preference Exploration by Monotonic Neural Network Ensemble), a novel approach to multi-objective optimization that learns the decision maker’s preferences, leveraging monoticity of the underlying utilities. The experimental results demonstrate the effectiveness of this method across a range of test problems. BOPE-MoNNE consistently outperformed other methods, including BOPE-GP (Lin et al., 2022) and PBO (Astudillo et al., 2023), across various test problems. A key strength of the proposed approach lies in its ability to incorporate monotonicity of the utilities through our neural network design, addressing a crucial property of utility functions that is often overlooked in existing approaches.
Looking ahead, there are several promising directions for future work. This paper identified challenges faced by HMC-trained BNNs in preference learning, particularly due to the scale invariance of preference learning and the independent prior distributions. An important area of exploration is how to effectively utilize HMC to train BNNs given pairwise comparison, as an alternative to ensemble-based approaches. Additionally, we are interested in extending the work to group decision-making scenarios, where multiple decision makers are involved. Preference-based multi-objective optimization has a wide range of applications, from science over engineering to business problems that we expect to benefit from the advancements presented in this work.
@@ A.1 Algorithm
Algorithm 1 MoNNE as the utility model in BOPE
1:Initial observation $\mathcal{D}_{0}$, initial pairwise comparison $\mathcal{P}_{0}$, and $\mathcal{R}_{0}$
@@
8:  Update $\mathcal{D}_{t+1}=\mathcal{D}_{t}\cup\{x_{t+1},y_{t+1}\}$
@@
10:  Find $\{y_{1,t+1},y_{2,t+1}\}\in\arg\max_{\{y_{1,i},y_{2,i}\}}\{\text{IEUBO}(f_{\text{true}}(x_{1,i}),f_{\text{true}}(x_{2,i})|g)\},$
11:    where $\{y_{1,i},y_{2,i}\}\in\{y_{i}|(x_{i},y_{i})\in D_{n}\}$
12:  Query preference:
13:    $p_{t+1}=2\cdot\mathbbm{1}_{g_{\text{true}}(y_{1,t+1})-g_{\text{true}}(y_{2,t+1})+\epsilon>0}-1$, $\epsilon\sim\mathcal{N}(0,\sigma_{\text{noise}}^{2})$
14:  Update $\mathcal{P}_{t+1}=\mathcal{P}_{t}\cup\{y_{1,t+1},y_{2,t+1}\}$ and $\mathcal{R}_{t+1}=\mathcal{R}_{t}\cup\{p_{t+1}\}$
15:end for
@@ A.3 HMC-trained BNN Given Pairwise Comparisons
A.4 Model Visualization Given Different Number of Pairwise Comparisons
@@ A.5.1 Noise Model
In this paper, we consider noise in the utility values, which can lead to errors in the decision maker’s choices. Another type of noise can be introduced through the Bradley-Terry model.
@@
Figure 11: Bradley-Terry and Gaussian noise models exhibit similar behavior
@@ A.5.3 Benchmark Algorithms
    qNParEGO [Daulton et al., 2020]: A multi-objective optimization algorithm that operates without learning decision maker (DM) preferences. Multi-objective optimization algorithms can be considered as strategies where the experimenter tries to find promising solutions on the Pareto front and asks the DM for their preference at the end.
@@
    PBO [Lin et al., 2022, Astudillo et al., 2023]: An approach that learns DM preferences directly in the solution space $\mathcal{X}$. It constructs a model $g(x)$ to predict the DM preference for a solution $x$. The acquisition functions are defined as $\operatorname{qNEIUU}\left(x_{1:q}\right)=\frac{1}{M}\sum_{j=1}^{M}\left\{\max g^{j}(x_{1:q})-\max g^{j}(X_{n}\right))\}^{+}$ and $\operatorname{EUBO}\left(x_{1,t},x_{2,t}\right)=\frac{1}{M}\sum_{j=1}^{M}\max\left\{g^{j}(x_{1,i}),g^{j}(x_{2,i})\right\}$. It is important to note that we still require an output surrogate model $f$ as DM can only compare the output pair rather than the input pairs. As the true output function is unknown, we use the posterior samples of $(f(x_{1,i}),f(x_{2,i}))$ as output pairs. As mentioned in the literature review, PBO is different from BOPE as their information sources are different. In the BOPE problem setting, DMs compare solutions based on their objective function values, while in PBO, comparisons are based on their design variables, making PBO not directly applicable. Therefore, in our benchmark PBO algorithm implementation, we construct a surrogate model of the output function and use output samples from pairs selected by PBO to elicit DM preferences in the Preference Exploration stage. In the Experimentation stage, we use the Expected Improvement acquisition function to pick a solution to be evaluated. Its result is shown in Figure 3. However, this setting may show the DM any arbitrary output pairs instead of only observed ones. Another way to make PBO applicable is to do two experimentation evaluations combined and then one preference exploration. This approach is necessary for PBO because a DM cannot directly compare two different inputs; instead, we must first obtain the outputs from both methods and subsequently elicit the DM’s preferences between these outputs. We conducted experiments using also this more equitable framework (one preference exploration and two experimentation evaluations in every iteration). Results presented in Figure 12 show that the order of performance of PBO, BOPE-GP and BOPE-MONNE remains the same.
@@ A.6 Activation Function Comparison
During the empirical investigation, we observed frequent numerical instabilities when using Tanh and ReLU activation functions, which led us to focus our analysis on three alternatives: Sigmoid, Swish, and Leaky ReLU. These functions provide an instructive comparison between smooth (Sigmoid, Swish) and non-smooth (Leaky ReLU) activation functions. For Leaky ReLU, the negative slope parameter is set to 0.25. Figure 16 presents the performance comparison of MoNNE in these activation functions. While all three activation functions demonstrated comparable performance, Swish exhibited marginally superior results.
Although Leaky ReLU achieved performance comparable to that of Swish in the previous experiments, it exhibited limitations in modeling complex utility functions. Figure 15 illustrates these limitations by displaying the base models of MoNNE with different activation functions, trained on 30 randomly sampled pairs from the constraint surface $y_{1}+y_{2}=10$. We examine the same complex function previously analyzed in Figure 2(d) of Section 3, plotting the model values along the constraint surface $y_{1}+y_{2}=10$. The visualization reveals that Leaky ReLU struggles to capture the intricate patterns of the underlying function.
Figure 15: Given the shallow architecture of the neural network, Leaky ReLU activation functions may not be ideal for modeling complex underlying functions.
Figure 16: While all three activation functions demonstrated comparable performance, Swish exhibited marginally superior results.
@@ A.7 Layer Number Comparison
We evaluate three different network depths across five test problems, as shown in Figure 18. While the optimal number of layers varies across problems, we observe that the 3-layer MoNNE consistently achieves intermediate performance. Notably, the parameter count increases substantially with network depth: the 2-layer MoNNE contains hundreds of parameters, the 3-layer variant introduces approximately 1,000 additional parameters, and the 4-layer architecture further adds 10,000 parameters. Moreover, the 4-layer MoNNE exhibits significantly longer training times compared to shallower architectures. Based on this trade-off between performance and computational efficiency, we adopt the 3-layer MoNNE architecture for all subsequent experiments in this paper.
@@ A.9 Influence of Utility Noise
Figure 21: Comparison of Preference Errors in BOPE-GP and BOPE-MoNNE. BOPE-MoNNE generally has more errors. Notably, in subfigures (c) and (e), the error count for BOPE-MoNNE plateaus due to several runs reaching the error threshold and stopping.
We conducted experiments with different noise levels ($\sigma_{noise}=0$, $\sigma_{noise}=0.1$ and $\sigma_{noise}=0.5$) to investigate the impact of noise on both the GP models and the MoNNE model, with results shown in Figure 22. Our analysis reveals that BOPE-MoNNE consistently outperforms BOPE-GP across all noise levels. This performance difference is particularly noteworthy in the VehicleSafety problem, where a noise level of $\sigma_{\text{noise}}=0.5$ represents significant disruption, given that the output range of this problem is approximately $0.6$. Under these conditions, pairwise comparisons become nearly random. While BOPE-GP performance deteriorates substantially in this high-noise scenario, BOPE-MoNNE maintains reasonable effectiveness by preserving the underlying monotonicity constraints of the problem.
@@ A.10 Derivation of Expected Maximum of Two Normal Distributions
$\mathbb{E}[\max(X,Y)]=\mathbb{E}[X|X>Y]\mathbb{P}(X>Y)+\mathbb{E}[Y|Y>X]\mathbb{P}(Y>X)$
@@
$Z\sim\mathcal{N}(\mu_{1}-\mu_{2},\sigma_{1}^{2}+\sigma_{2}^{2})$
@@
$\mathbb{P}(X>Y)$ $=\mathbb{P}(Z>0)=\Phi(\alpha)$
$\mathbb{P}(Y>X)$ $=\mathbb{P}(Z<0)=\Phi(-\alpha)=1-\Phi(\alpha)$
@@
$\begin{pmatrix}X\\ Z\end{pmatrix}\sim\mathcal{N}\left(\begin{pmatrix}\mu_{1}\\ \mu_{1}-\mu_{2}\end{pmatrix},\begin{pmatrix}\sigma_{1}^{2}&\sigma_{1}^{2}\\ \sigma_{1}^{2}&\sigma_{1}^{2}+\sigma_{2}^{2}\end{pmatrix}\right)$
@@
$=\mu_{1}+\frac{\sigma_{1}^{2}}{\sigma_{3}}\frac{\phi(\alpha)}{\Phi(\alpha)}$
@@
$\mathbb{E}[Y|Y>X]=\mu_{2}+\frac{\sigma_{2}^{2}}{\sigma_{3}}\frac{\phi(-\alpha)}{\Phi(-\alpha)}$
@@
$\mathbb{E}[\max(X,Y)]=\left(\mu_{1}+\frac{\sigma_{1}^{2}}{\sigma_{3}}\frac{\phi(\alpha)}{\Phi(\alpha)}\right)\Phi(\alpha)+\left(\mu_{2}+\frac{\sigma_{2}^{2}}{\sigma}\frac{\phi(-\alpha)}{\Phi(-\alpha)}\right)\Phi(-\alpha)$
@@
$\boxed{\mathbb{E}[\max(X,Y)]=\mu_{1}\Phi(\alpha)+\mu_{2}\Phi(-\alpha)+\sigma_{3}\phi(\alpha)}$

### v3

@@ 6 Conclusion
This paper introduces BOPE-MoNNE (Bayesian Optimization with Preference Exploration using a Monotonic Neural Network Ensemble), a novel approach to multi-objective optimization that learns the decision maker’s preferences, leveraging monotonicity of the underlying utilities. The experimental results demonstrate the effectiveness of this method across a range of test problems. BOPE-MoNNE consistently outperforms other methods, including BOPE-GP (Lin et al., 2022) and PBO (Astudillo et al., 2023), across various test problems. A key strength of the proposed approach lies in its ability to incorporate monotonicity of utilities through our neural network design, addressing a crucial property of utility functions that is often overlooked in existing approaches.
Looking ahead, there are several promising directions for future work. This paper identified challenges faced by HMC-trained BNNs in preference learning, particularly due to the scale invariance of preference learning and the independent prior distributions. An important area of exploration is how to effectively utilize HMC to train BNNs given pairwise comparisons, as an alternative to ensemble-based approaches. Additionally, we are interested in extending the work to group decision-making scenarios, where multiple decision makers are involved. Preference-based multi-objective optimization has a wide range of applications, from science over engineering to business problems that we expect to benefit from the advancements presented in this work.
@@ A.1 Algorithm
Algorithm 1 BOPE-MoNNE
1:Initial observation $\mathcal{D}_{0}$, initial pairwise comparisons $\mathcal{P}_{0}$, and $\mathcal{R}_{0}$
@@
8:  Update $\mathcal{D}_{t+1}=\mathcal{D}_{t}\cup\{(x_{t+1},y_{t+1})\}$
@@
10:  Find $(y_{1,t+1},y_{2,t+1})\in\arg\max_{\{y_{1,i},y_{2,i}\}}\{\text{IEUBO}(y_{1,i},y_{2,i}|g)\},$
11:    where $y_{1,i},y_{2,i}\in\{y_{i}|(x_{i},y_{i})\in\mathcal{D}_{n}\}$
12:  Query preference: $p_{t+1}=2\cdot\mathbbm{1}_{g_{\text{true}}(y_{1,t+1})-g_{\text{true}}(y_{2,t+1})+\epsilon>0}-1$, $\epsilon\sim\mathcal{N}(0,\sigma_{\text{noise}}^{2})$
13:  Update $\mathcal{P}_{t+1}=\mathcal{P}_{t}\cup\{(y_{1,t+1},y_{2,t+1})\}$ and $\mathcal{R}_{t+1}=\mathcal{R}_{t}\cup\{p_{t+1}\}$
14:end for
@@ A.3 HMC-trained BNN Given Pairwise Comparisons
A.4 Model Visualization Given Different Numbers of Pairwise Comparisons
@@ A.5.1 Noise Model
In this paper, we consider noise in utility values, which can lead to errors in the DM’s choices. Another type of noise can be introduced through the Bradley-Terry model.
@@
Figure 11: Bradley-Terry and Gaussian noise models exhibit similar behaviour.
@@ A.5.3 Benchmark Algorithms
    qNParEGO [Daulton et al., 2020]: A multi-objective optimization algorithm that operates without learning DM preferences. Multi-objective optimization algorithms can be considered as strategies where the experimenter tries to find promising solutions on the Pareto front and asks the DM for their preference at the end.
@@
    PBO [Lin et al., 2022, Astudillo et al., 2023]: An approach that learns DM preferences directly in the solution space $\mathcal{X}$. It constructs a model $g(x)$ to predict the DM preference for a solution $x$. The acquisition functions are defined as $\operatorname{qNEIUU}\left(x_{1:q}\right)=\frac{1}{M}\sum_{j=1}^{M}\left\{\max g^{j}(x_{1:q})-\max g^{j}(X_{n}\right))\}^{+}$ and $\operatorname{EUBO}\left(x_{1,t},x_{2,t}\right)=\frac{1}{M}\sum_{j=1}^{M}\max\left\{g^{j}(x_{1,i}),g^{j}(x_{2,i})\right\}$. It is important to note that we still require an output surrogate model $f$ as DM can only compare the output pair rather than the input pairs. As the true output function is unknown, we use the posterior samples of $(f(x_{1,i}),f(x_{2,i}))$ as output pairs. As mentioned in the literature review, PBO is different from BOPE as their information sources are different. In the BOPE problem setting, DMs compare solutions based on their objective function values, while in PBO, comparisons are based on their design variables, making PBO not directly applicable. Therefore, in our benchmark PBO algorithm implementation, we construct a surrogate model of the output function and use output samples from pairs selected by PBO to elicit DM preferences in the Preference Exploration stage. In the Experimentation stage, we use the Expected Improvement acquisition function to pick a solution to be evaluated. Its result is shown in Figure 3. However, this setting may show the DM any arbitrary output pairs instead of only observed ones. Another way to make PBO applicable is to do two experimentation evaluations combined and then one preference exploration. This approach is necessary for PBO because a DM cannot directly compare two different inputs; instead, we must first obtain the outputs from both methods and subsequently elicit the DM’s preferences between these outputs. We conducted experiments using this more equitable framework (one preference exploration and two experimentation evaluations in every iteration). Results presented in Figure 12 show that the order of performance of PBO, BOPE-GP and BOPE-MONNE remains the same.
@@ A.6 Activation Function Comparison
During the empirical investigation, we observed frequent numerical instabilities when using Tanh and ReLU activation functions, which led us to focus our analysis on three alternatives: Sigmoid, Swish, and Leaky ReLU. These functions provide an instructive comparison between smooth (Sigmoid, Swish) and non-smooth (Leaky ReLU) activation functions. For Leaky ReLU, the negative slope parameter is set to 0.25. Figure 16 presents the performance comparison of MoNNE in these activation functions. While all three activation functions demonstrate comparable performance, Swish exhibits marginally superior results.
Although Leaky ReLU achieves performance comparable to that of Swish in the previous experiments, it exhibits limitations in modelling complex utility functions. Figure 15 illustrates these limitations by displaying the base models of MoNNE with different activation functions, trained on 30 randomly sampled pairs from the constraint surface $y_{1}+y_{2}=10$. We examine the same complex function previously analyzed in Figure 2(d) of Section 3, plotting the model values along the constraint surface $y_{1}+y_{2}=10$. The visualization reveals that Leaky ReLU struggles to capture the intricate patterns of the underlying function.
Figure 15: Given the shallow architecture of the neural network, Leaky ReLU activation functions may not be ideal for modelling complex underlying functions.
Figure 16: While all three activation functions demonstrate comparable performance, Swish exhibits marginally superior results.
@@ A.7 Layer Number Comparison
We evaluate three different network depths across six test problems, as shown in Figure 18. While the optimal number of layers varies across problems, we observe that the 3-layer MoNNE consistently achieves intermediate performance. Notably, the parameter count increases substantially with network depth: the 2-layer MoNNE contains hundreds of parameters, the 3-layer variant introduces approximately 1,000 additional parameters, and the 4-layer architecture further adds 10,000 parameters. Moreover, the 4-layer MoNNE exhibits significantly longer training times compared to shallower architectures. Based on this trade-off between performance and computational efficiency, we adopt the 3-layer MoNNE architecture for all subsequent experiments in this paper.
@@ A.9 Influence of Utility Noise
Figure 21: Comparison of Preference Errors in BOPE-GP and BOPE-MoNNE. BOPE-MoNNE generally has more errors. Notably, in the VehicleSafety problem and the CarCabDesign problem, the error count for BOPE-MoNNE plateaus due to several runs reaching the error threshold and stopping.
We conducted experiments with different noise levels ($\sigma_{\text{noise}}=0$, $\sigma_{\text{noise}}=0.1$ and $\sigma_{\text{noise}}=0.5$) to investigate the impact of noise on both the GP model and the MoNNE model, with results shown in Figure 22. Our analysis reveals that BOPE-MoNNE consistently outperforms BOPE-GP across all noise levels. This performance difference is particularly noteworthy in the VehicleSafety problem, where a noise level of $\sigma_{\text{noise}}=0.5$ represents significant disruption, given that the output range of this problem is approximately $0.6$. Under these conditions, pairwise comparisons become nearly random. While BOPE-GP performance deteriorates substantially in this high-noise scenario, BOPE-MoNNE maintains reasonable effectiveness by preserving the underlying monotonicity constraints of the problem.
@@ A.10 Derivation of Expected Maximum of Two Normal Distributions
$\mathbb{E}[\max(X,Y)]=\mathbb{E}[X|X>Y]\mathbb{P}(X>Y)+\mathbb{E}[Y|Y>X]\mathbb{P}(Y>X).$
@@
$Z\sim\mathcal{N}(\mu_{1}-\mu_{2},\sigma_{1}^{2}+\sigma_{2}^{2}).$
@@
$\mathbb{P}(X>Y)$ $=\mathbb{P}(Z>0)=\Phi(\alpha),$
$\mathbb{P}(Y>X)$ $=\mathbb{P}(Z<0)=\Phi(-\alpha)=1-\Phi(\alpha),$
@@
$\begin{pmatrix}X\\ Z\end{pmatrix}\sim\mathcal{N}\left(\begin{pmatrix}\mu_{1}\\ \mu_{1}-\mu_{2}\end{pmatrix},\begin{pmatrix}\sigma_{1}^{2}&\sigma_{1}^{2}\\ \sigma_{1}^{2}&\sigma_{1}^{2}+\sigma_{2}^{2}\end{pmatrix}\right)$.
@@
$=\mu_{1}+\frac{\sigma_{1}^{2}}{\sigma_{3}}\frac{\phi(\alpha)}{\Phi(\alpha)}.$
@@
$\mathbb{E}[Y|Y>X]=\mu_{2}+\frac{\sigma_{2}^{2}}{\sigma_{3}}\frac{\phi(-\alpha)}{\Phi(-\alpha)}.$
@@
$\mathbb{E}[\max(X,Y)]=\left(\mu_{1}+\frac{\sigma_{1}^{2}}{\sigma_{3}}\frac{\phi(\alpha)}{\Phi(\alpha)}\right)\Phi(\alpha)+\left(\mu_{2}+\frac{\sigma_{2}^{2}}{\sigma}\frac{\phi(-\alpha)}{\Phi(-\alpha)}\right)\Phi(-\alpha).$
@@
$\boxed{\mathbb{E}[\max(X,Y)]=\mu_{1}\Phi(\alpha)+\mu_{2}\Phi(-\alpha)+\sigma_{3}\phi(\alpha)}.$